# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v3

@@ Abstract.
We investigate youth visions for ideal remote social interactions, drawing on co-design interviews with 23 participants (aged 15-24) experienced with 3D gaming environments. Using a Fictional Inquiry (FI) method set in the Harry Potter universe, this research reveals that young people desire social media that functions more like immersive, navigable shared social spaces. Across these interviews, participants identified six key priorities for meaningful social connection over social media: intuitive social navigation, shared collaborative experiences, communal environments fostering close relationships, flexible self-presentation, intentional engagement, and playful social mechanics. We introduce the spatial integrity framework, a set of four interrelated design principles: spatial presence, spatial composition, spatial configuration, and spatial depth. Together, these principles outline how online spaces can be designed to feel more like meaningful environments, spaces where relationships can grow through shared presence, movement, and intentional interaction. Participants also described the FI process itself as meaningful, not only for generating new ideas but for empowering them to imagine and shape the future of social media.
@@ 1. Introduction
Meaningful social connections are critical for young people’s development and well-being (Hawkley and Cacioppo, 2010), and social media has great potential to support these. At the individual level, richer and healthier interactions support youth well-being and identity development (Kim et al., 2024c). At the societal level, platforms act as critical infrastructures for sustaining friendships and community ties, making their design choices consequential for social capital (Phua et al., 2017a; Ellison et al., 2007). In fact, many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n. d.]).
However, social media platforms increasingly shift away from facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025). While designers initially created these platforms around sharing personal updates with friends, they now emphasize entertainment, news, and engaging larger, less intimate audiences. Meta’s data illustrates this shift: between 2023 and 2025, time spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025). Mark Zuckerberg described this change as aligned with the company’s focus on “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
This shift raises concerns about meaningful online relationships, particularly for young people who have grown up with social media as central to their social lives. Most young people engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023). Despite high engagement, many report dissatisfaction with their social interactions, not because they want reduced usage, but because they expect platforms to better support their connections (Kim et al., 2024c). Current designs prioritize passive content consumption over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
Given social connections’ importance for young people and social media’s vital role in facilitating them, we must explore new design possibilities. Young people frequently express dissatisfaction with current experiences, revealing a significant gap between their social needs and existing platform interactions. Mainstream platforms employ standardized design patterns that constrain creative possibilities and foster pessimism among users about social media’s potential. This creates a compelling need for innovative designs that complement existing platforms by emphasizing meaningful interpersonal connections. Therefore, we ask the following research question: What types of social interactions might current social media designs be overlooking or inadequately supporting?
To address this question, we conducted co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) with 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n. d.]) and Roblox (Roblox Corporation, [n. d.]). Participants collaboratively imagined a social media platform designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This imaginative context allowed participants to step away from existing platform constraints and envision more varied and meaningful social interactions. Since researchers recognize 3D gaming environments as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated participants could leverage gaming insights to propose novel social media interactions. Specifically, our research addresses: What types of interactions do young people familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
Our findings reveal that young people with 3D gaming experience envision ideal social media platforms as immersive, navigable environments that designers should create to feel like neighborhoods, community centers, and shared social spaces. Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming. These visions highlight young people’s desire for social media that feels more relational, meaningful, and personally rewarding.
We contribute empirical findings on young people’s perceptions of current social media platforms’ shortcomings and their desired experiences. We propose the concept of spatial integrity for social interactions in online spaces, encompassing four dimensions of spatial affordances that facilitate meaningful social connections: spatial presence (understanding the space and identifying who is present), spatial composition (the space’s meaning and significance), spatial configuration (the space’s organizational structure), and spatial depth (the extent of perceivable information within the space). We demonstrate that the FI method in social media design contexts prompts creative idea generation, enables participants to transcend current design patterns, and empowers them to envision social media’s positive potential, leaving them more hopeful about future possibilities.
@@ 2. Related Work
2.1. Meaningful Social Connection in Online Environments
2.1.1. Defining Meaningful Social Interaction
@@
Litt et al. (Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” Though “deeply subjective,” these interactions are distinguished by their impact “beyond the immediate interaction itself,” involving positive emotions, information exchange, or concrete benefits.
The study identified three key impact types: “emotional impact” (feelings, mood changes, or relationship strengthening), “informational impact” (knowledge, advice, or understanding exchange), and “tangible impact” (concrete benefits like goods or services). Non-meaningful interactions lacked any such impact.
Strong ties (friends and family), community ties, and shared activities were associated with meaningfulness, but acted as facilitators enabling impact rather than primary sources of meaning. Other attributes linked to meaningfulness included advance planning, memorialization (photos or videos), and synchronicity. The communication medium itself—online or offline—was less important than these other characteristics.
@@
We define meaningful social interaction as any interaction that participants would actively seek in an ideal scenario, reflecting their subjective meaningfulness. We expected these themes to emerge in participants’ descriptions of their experiences.
2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
Though social media platforms are primary connection tools for young people (Boyd, 2014a), many users report feeling lonely or disconnected despite widespread use (Twenge et al., 2019; Primack et al., 2017), highlighting a gap between intended purpose and actual experience.
@@ 2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
While social connection remains a primary motivation for social media use, mainstream platforms like Instagram (All, [n. d.]), YouTube (you, [n. d.]), and Discord (discord page, [n. d.]) are designed for diverse user motivations including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022). From a uses and gratifications perspective (Katz et al., 1973), users pursue varied goals from building social capital (Phua et al., 2017b) to media consumption (Menon, 2022). Users seeking social connection may encounter experiences prioritizing other engagement forms, leaving social needs unmet.
This gap underscores the need for a clearer understanding of meaningful social connections in online environments. While social media presents risks like social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b) alongside benefits like marginalized group support (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and identity development opportunities (Davis, 2012; Lee et al., 2022), social connection remains central for many users. For young people increasingly relying on digital platforms, we must move beyond treating connection as one competing goal toward a focused understanding of how platforms can meaningfully support social relationships.
2.2. Social Technology Design Explorations in HCI
2.2.1. Designing for Meaningful Social Connections
HCI research has extensively explored designs fostering meaningful social interactions online at both platform and feature levels. Platforms like BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) introduced unique affordances emphasizing spontaneity, ephemerality, and creative self-expression. This research identified ways to build relationships through features prioritizing authenticity and playful interaction. However, users report increasingly toxic interactions as platforms gain popularity, with design choices encouraging network expansion and compulsive behaviors like tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), detracting from meaningful interaction goals.
Feature-level efforts include designs encouraging effortful communication (Liu et al., 2021; Zhang et al., 2022) and genuine connections (Stepanova et al., 2022). However, social media experiences are holistic, involving a complex interplay of multiple design aspects and user behaviors. Individual features alone cannot fundamentally alter overarching social media narratives.
Beyond archetypal social media, recent work investigates spatial social technologies emphasizing immersive, embodied interactions. Social VR platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide virtual spaces and embodiment for co-presence and shared experiences, facilitating more personal and tangible interactions than 2D environments. Platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend gaming and social networking, enabling virtual environment navigation, avatar customization, and dynamic group interactions. However, these platforms are not generally recognized as conventional social media.
What youth envision as ideal social media at the landscape and feature levels remains unclear, especially considering emerging technologies like spatial computing platforms. Understanding these perspectives is crucial for reimagining social media’s future to meet users’ evolving needs.
2.2.2. Video Games as a Source of Meaningful Social Connection
Video games are increasingly recognized as valuable environments for meaningful social connection, particularly among youth. Large-scale surveys show games serve as major venues for friendship development and maintenance, especially for boys (Lenhart, 2015c). Teens play with known friends and make new ones through online gameplay (Lenhart, 2015a), with over half reporting meeting at least one friend online (Lenhart, 2015a, b). These interactions extend beyond gameplay, supporting ongoing relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
Research highlights gaming’s social and emotional benefits, including cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014). Certain designs promote prosocial outcomes by encouraging players to help others in-game and beyond (Gentile et al., 2009). Games offer deeper, eudaimonic experiences involving meaning-making, personal growth, and connection (Daneels et al., 2023). These findings suggest video games can foster meaningful social connections among youth and highlight the potential for game-like features in social media.
@@ 2.3. Recognizing and Overcoming Social Media Design Fixation
Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented design challenge. Though generating divergent ideas is critical for innovation, we often impose unnecessary restrictions. This self-limiting process confines design to localized problems rather than exploring broader, optimal solutions.
Social media design faces similar fixation (Sundaram, [n. d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020). Ad-based revenue models incentivize increased time on apps (Oth, [n. d.]). Platforms use compulsive strategies like friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)), creating distrustful user communities (Kim et al., 2025). Rather than proactively exploring new paradigms of designs and interactions, much recent work on social media design has disproportionately focused on reactively addressing existing harm (Kim et al., 2024b). The domination of mainstream platforms along with this reactive approach to design leaves users dissatisfied and reinforces perceptions that all social media experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
Analogical transfer offers one powerful method for overcoming design fixation (Novick, 1988; Buchanan, 1992). Therefore, in this paper, we attempt to break out of the design and cognitive fixation that assumes the future of social media is bleak by freeing participants from real-world constraints and beliefs and leveraging elements from fictional settings to serve as sources for analogical transfer.

### v4

# Social Media Should Feel Like Minecraft, Not Instagram: Youth Visions for Meaningful Social Connections through Fictional Inquiry

@@ Abstract.
We conducted co-design workshops with 23 participants (ages 15–24) to explore how youth envision ideal remote social connection. Using Fictional Inquiry (FI) within a Harry Potter-inspired narrative, we found that youth perceive a disconnect between platforms labeled “social media” (like Instagram) and those where they actually experience meaningful connections (like Minecraft or Discord). Participants envisioned an immersive platform prioritizing meaningful social connection through presence and immersion, natural self-expression, intuitive spatial navigation leveraging physical-world norms, and playful, low-stakes opportunities for friendship development. We synthesize these visions into six themes articulating relational needs that current platforms systematically marginalize. The FI method proved effective in generating innovative ideas while empowering youth by fostering hope and agency over social media’s future. Our findings challenge “doom” narratives by reframing social media’s harms as outcomes of specific design choices, demonstrating how design research can reopen space for imagining more supportive forms of mediated connection.
@@ 1. Introduction
Meaningful social connections are critical for young people’s development and well-being (Hawkley-2010-LonelinessMattersMechanisms-v). Early social computing researchers and users celebrated these technologies for their potential to build relationships that transcend physical boundaries (ellison2007benefits). Over time, however, that promise has narrowed. Between 2023 and 2025, time spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley-2025-InstagramFacebookProof-l). Current designs increasingly prioritize passive consumption over active interaction with close friends (Landesman-2024-IInstagram-j), and are associated with fatigue, mistrust, and perceived harm. Platforms once imagined as infrastructures for personal connection now make even small acts of self-disclosure—the basic currency of relationship development—feel high-stakes (Kim2025Trust-EnabledRegulationz) and anxiety-inducing (jaewon-privacy).
For youth who have grown up with these technologies as central to their social lives, this shift carries consequences beyond user experience. While young people report frustration, anxiety, and toxicity, they also describe moments of genuine joy and meaningful connection—and participation remains a practical necessity for staying connected with peers. Yet dominant policy, educational, and design responses frame youth engagement in terms of prevention and protection, emphasizing restriction over support for experimentation that might foster agency and resilience (Kim2025DesignChallengese; Kim2024EnvisioningPreventingn). Youth consistently report that measures such as outright bans leave them feeling unheard (Kim-2024-AustraliaBarred-y; TheLearningNetwork-2025-WhatTeensMedia-z; Weinstein2022-rh). Their relationships with social media are ambivalent rather than uniformly negative, but this complexity is rarely reflected in prevailing approaches to governance or design.
This disconnect is reinforced by the broader convergence of social media platforms themselves (Sundaram-Other-SocialMediaSame-g; Pardes-2020-SocialMediaSame-p; Pitt-2021-SocialMediaSame-s; Nielsen-2023-ComparingPlatformizationAnalysis-m; VanDyke-2020-GrowingSimilaritiesPlatforms-m). Despite surface-level differences, mainstream platforms increasingly share a common paradigm organized around feeds, mobile interfaces, and vertical profiles, crowding out alternative models of mediated connection. As a result, much design research now operates in a reactive mode. The guiding question has shifted from what meaningful online connection might look like to how people can be protected from existing platforms. This framing treats the dominant paradigm as fixed, focusing on mitigation rather than reimagination. Although generative work exists within HCI and related fields, it is often overshadowed by protectionist narratives that position youth as vulnerable users in need of safeguards, prioritizing safety over empowerment and incremental repair over structural alternatives (Kim2025DesignChallengese; Kim2024EnvisioningPreventingn).
These constraints shape not only policy and design discourse but also participatory research itself. When young people are asked to redesign social media, the prompt arrives burdened with existing controversies around addiction, mental health, and toxicity. Participants often respond defensively, reproducing adult-driven narratives through social desirability bias even when their lived experiences are more nuanced (Davis2019EverythingsRelationshipsg). Without language or precedents beyond existing platforms, participants tend to remix familiar features—a cleaner feed, better moderation, fewer metrics—rather than imagining fundamentally different forms of social interaction. The term “social media” itself anchors imagination to platforms like Instagram or TikTok, constraining what can be envisioned before design even begins. As a result, participatory design risks generating refinements when its prompts and methods keep participants inside the same conceptual frame. To move beyond these constraints, we ask:
  • RQ: What forms of social interaction do youth envision when freed from current platform constraints?
Rather than beginning from existing systems or problems, we focus on the relationships, interactions, and forms of presence young people desire, and on the design considerations that emerge from these visions.
To address this question, we conducted co-design interviews using the Fictional Inquiry (FI) method (IversenDindler-2007-FictionalInquiry--designSpace-m), a speculative approach that invites participants to imagine within a fictional scenario removed from real-world constraints. Twenty-three participants aged 15–24 imagined how students at Hogwarts—the fictional school from Harry Potter(harrypotterfanclub)—might connect with friends beyond their campus using any magical powers they could envision. This framing served three purposes. First, it functioned as a defamiliarization strategy, breaking the conceptual lock of existing platforms by removing both technical limitations and loaded terminology, starting from desire rather than constraint. Second, it centered meaningful social connection as the explicit design goal, reflecting youth developmental needs and interests documented in prior research (Davis2012-bq). Third, it created a space of play in which youth are natural experts, enabling them to express values, social logics, and relational priorities difficult to articulate within dominant narratives about social media. Participants were not asked to redesign social media; they were asked how wizards might stay close across distance.
Our findings reveal six themes characterizing how youth envision meaningful remote social interaction: intuitive social navigation modeled on real-world movement and presence; shared experiences through collaborative activities and casual co-presence; communal environments that prioritize close relationships and low-pressure interaction; flexible self-presentation through avatars and contextual identity management; intentional engagement that replaces passive scrolling with purposeful navigation; and playful, low-stakes social mechanics. Within these themes, participants described concrete interactions—sitting together in a virtual study room with Lo-Fi music, leaving a small gift at a friend’s door without initiating a full conversation, or using secret knocks to grant trusted friends access to private spaces. Although participants developed their visions independently, their accounts of how these imagined systems would feel converged strongly, emphasizing connection, engagement, fulfillment, and agency. These patterns suggest shared relational needs not well supported by current platforms. We also observed that the FI process itself shifted participants from reactive problem-fixing toward aspirational imagining, with many reporting increased hope and excitement about what mediated connection could become.
This paper makes three contributions. First, we offer a methodological account of FI as a research-through-design approach for reimagining meaningful remote social connection. By suspending constraints and removing the term “social media” itself, the fictional framing enabled youth—a population often positioned as subjects to protect rather than agents of imagination—to reason from felt experience rather than inherited platform templates. Second, we present an empirical account of the interaction concepts, affordances, and social mechanisms young people proposed when freed from feed-based paradigms. Third, we synthesize these concepts into six themes that articulate underlying relational needs and design directions, highlighting what modern social media systematically marginalizes. Together, these contributions challenge “doom” narratives around social media, reframing them as outcomes of specific design choices and demonstrating how design research can reopen space for imagining more supportive forms of mediated connection.
@@ 2. Related Work
2.1. Beyond Moral Panic: From Panic Narratives to Design Opportunity
Widespread societal concerns exist about social media’s impact on youth well-being (OtherOther-2023-SocialMediaAdvisory-u). Rather than addressing these constructively, public discourse has framed social media as a societal evil (Wike-2022-2ViewsSociety-y), particularly in the U.S., where it is blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt-2023-SocialMediaEvidence-c). This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
However, treating social media as inherently harmful, or youth as best served by removal, collapses a more complex reality. Despite public concern about social media’s role in youth mental health, causal evidence remains inconclusive, and youth report a mix of positive and negative experiences (Vogels-2023-TeensSocialSurveys-n). Fear-based policies and alarmist narratives can themselves cause harm, amplifying anxiety, mistrust, and helplessness among youth (kim2024privacysocialnormsystematically; lee2024social). Recognizing nuance does not deny that serious harms occur, including high-risk vulnerabilities (e.g., sexual exploitation, grooming, and synthetic media threats), nor ignore structural forces like platform power asymmetries and algorithmic manipulation. Instead, it motivates a complementary design question: what kinds of online social environments could better support youth relationships while accounting for both everyday experiences and real risks?
2.2. Meaningful Social Connection Online, and Why Mainstream Platforms Miss It
What counts as meaningful connection
@@ 2.2. Meaningful Social Connection Online, and Why Mainstream Platforms Miss It
Litt et al. (Litt-2020-WhatMeaningfulSurvey-h) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” Though “deeply subjective,” these interactions are distinguished by their impact “beyond the immediate interaction itself.” They identify three key impact types: “emotional impact” (feelings, mood changes, or relationship strengthening), “informational impact” (knowledge, advice, or understanding exchange), and “tangible impact” (concrete benefits like goods or services). Strong ties, community ties, and shared activities facilitate meaningfulness, while attributes like advance planning, memorialization, and synchronicity often support impact.
@@
Building on this work, we treat meaningful social interaction as any interaction participants would actively seek in an ideal scenario, reflecting their subjective sense of meaningfulness.
Connection-seeking on platforms optimized for other gratifications
@@
Though social media platforms are primary connection tools for young people (BoydBoyd-2014-ComplicatedSocialTeens-j), many users report feeling lonely or disconnected despite widespread use (twenge2019less; primack2017social), highlighting a gap between social media’s promise and lived experience. Mainstream platforms like Instagram (instagram), YouTube (youtube), and Discord (discord) serve diverse user motivations including entertainment, information seeking, and self-promotion (whiting2013people; Menon-2022-UsesGratificationsInstagram-x). From a uses and gratifications perspective (GurevitchKatz-1973-UsesGratificationsResearch-y), users pursue varied goals from building social capital (phua_uses_2017) to media consumption (Menon-2022-UsesGratificationsInstagram-x). Users who primarily seek meaningful connection may thus encounter experiences that prioritize other forms of engagement, leaving social needs unmet.
HCI and CSCW work has explored designs supporting more meaningful interaction, including efforts encouraging effortful communication (LiuFannie2021SOUt; LiuZhang-2022-AuggieEncouragingExperiences-n) and reflective or value-oriented social interaction (Stepanova2022-vh). More broadly, recent work articulates design spaces and system-level alternatives beyond dominant templates (e.g., form factors, interaction conventions, and social architectures) (zhang2024form). Yet if connection is so central, why do mainstream platforms consistently fail to prioritize it? Part of the answer lies in structural forces constraining both platform incentives and design imagination.
@@ 2.3. Recognizing and Overcoming Social Media Design Fixation
Design fixation (jansson1991design) describes the tendency to remain constrained by conventional thinking or existing solutions, even when divergent exploration is needed. Social media design faces similar fixation (Sundaram-Other-SocialMediaSame-g; Pardes-2020-SocialMediaSame-p; Pitt-2021-SocialMediaSame-s; Nielsen-2023-ComparingPlatformizationAnalysis-m; VanDyke-2020-GrowingSimilaritiesPlatforms-m). Ad-based revenue models incentivize increased time on apps (Other-Other-PlatformCapitalism-f), shaping platforms toward engagement metrics and features that prolong attention. Platforms deploy compulsive strategies like friend recommendations (HinikerKim-2024-SharingDesign-x) and engagement metrics (e.g., Snap scores (rozgonjuk2021comparing)), which erode trust and foster distrustful user communities (Kim2025Trust-EnabledRegulationz). In parallel, design efforts are often dominated by reactive harm mitigation rather than proactive exploration of new approaches (Kim2024EnvisioningPreventingn). Together, these forces help explain why users experience mainstream platforms as converging toward similar, feed-driven patterns, even when those patterns do not align with meaningful connection goals.
These constraints motivate methods that deliberately create distance from current systems, enabling participants to articulate what they would want social media to optimize for without starting from existing platform assumptions.
2.4. Alternative Paradigms and Methods for Imagining Beyond Platform Templates
Sociality in games and multi-space online communities
Video games are increasingly recognized as valuable environments for meaningful social connection, particularly among youth. Large-scale surveys show games serve as major venues for friendship development and maintenance (Lenhart-2015-Chapter3Boys-v), with teens playing with known friends and making new ones through online gameplay (Lenhart-2015-Chapter1Friendships-i). CSCW research shows that gaming communities often extend across multiple spaces, requiring coordination and group management across platforms and contexts (toombs2022we). These findings suggest gaming-influenced sociality can illuminate how people sustain relationships through shared activities, low-stakes co-presence, and community infrastructure, while raising questions about what translates and who may be excluded by more spatial or higher-cognitive-load interaction styles.
Spatial social technologies and place-based interaction
Beyond archetypal social media, HCI and CSCW have examined spatial social technologies emphasizing immersive and embodied interaction. Social VR and virtual worlds support co-presence, shared experiences, and identity play through avatars and navigable spaces (Zamanifard-2023-SurpriseBirthdayDistance-n; Freeman-2024-MyAudiences-u; Freeman-2020-MyBodyReality-l; Zamanifard-2019-TogethernessCraveRelationships-s; deighan2023social; rzeszewski2024social). Platforms like Gather.town (duarte2023experience; tu2022meetings) and VRChat provide place-like environments where spatial layout, navigation, and ambient co-presence shape interaction rhythms. A longer history of virtual worlds (e.g., Second Life) highlights that spatiality can enable rich social experience while introducing familiar challenges around governance, safety, and uneven participation. We draw on these literatures as precedents for treating online social spaces as places, where constructs from place theory, environmental psychology, and presence research (e.g., propinquity, wayfinding, place attachment, social presence, and co-presence) can help interpret how spatial organization relates to social connection.
Youth-centered design and fiction-based methods for escaping fixation
Youth-focused HCI has long argued for engaging young people as stakeholders and co-designers rather than solely as subjects of protection. HCI has also developed fiction-based and speculative methods that create distance from present-day systems, making it easier to reason about values, tradeoffs, and alternative social arrangements. Related approaches include design fiction, speculative design, and pastiche scenarios, which use crafted narratives and artifacts to surface new possibilities and critique existing assumptions (blythe2006pastiche; dunneraby2013speculative). Prior HCI work has drawn explicitly on fictional worlds, including Harry Potter, to design for social values and family coordination (brown2007locating).
Building on this methods lineage, we use FI to structure analogical transfer through a shared fictional narrative; unlike design fiction or pastiche scenarios that often center researcher-authored speculative artifacts, FI treats a well-known fictional world as a common design constraint and supports participant-led world-building of social architectures and norms. This approach is especially relevant for domains like social media, where incentive structures and design fixation can make it difficult to surface what youth might find meaningful outside current assumptions.